OpenStudy (anonymous):

If a running system has a total change in the heat of 315 joules, and it is running at a temperature of 350 K, what is the entropy? 1.54 joules/kelvin 0.847 joules/kelvin 1.22 joules/kelvin 0.900 joules/kelvin 0.762 joules/kelvin

OpenStudy (anonymous):

Entropy change = dQ/T = 315/350 = 0.900 J/oK